OpenStudy (anonymous):

. Replace ? with =, >, or < to make the statement true. 18 ÷ 6 + 3 ? 6 + 12 ÷ 3 A. = B. < C. >

OpenStudy (anonymous):

@ElShortie100

OpenStudy (anonymous):

@paigeRG

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

18/6=3 3+3= 6 6+12=18 18/3=6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

sooo they would be =

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

anyone correct me if im wrong

OpenStudy (anonymous):

@paigeRG didn't follow the correct order of operations. sorry. division first on the right then add

OpenStudy (anonymous):

okay so my answer is wrong?

OpenStudy (anonymous):

so its actually < not =

OpenStudy (anonymous):

yeah... the second one should be 12/3 = 4 and then 6+4 = 10

OpenStudy (anonymous):

ahhh I see well now I feel bad... but I also learned something....

OpenStudy (anonymous):

we have the rule BODMAS means Brackets Of Division Multiplication Addition Subtraction; means first divide then multiply then add then subtract  so answer is BBBBBBBBB